OpenStudy (anonymous):

use the quadratic formula to find the zeros of f. f(x) = -2x2 - 9x + 18

OpenStudy (anonymous):

just to clarify thats -2x^2

OpenStudy (anonymous):

-2x^2-9x+18=0 (-2x+3)(x+6)=0 x=3/2 and x=-6